given that p and q are real & that 1+2i are aroot of the equation z^2+(p+5i)z +q(2-i) determine the value of p& q and the other root of the equation
1
Expert's answer
2021-08-18T11:53:42-0400

Given that "p" and "q" are real and that "1+2i" are a root of the equation "z^2+(p+5i)z +q(2-i)" let us determine the value of "p" and "q" and the other root of the equation.

Let "a+bi" be the other root of the equation. Then


"z^2+(p+5i)z +q(2-i)=(z-1-2i)(z-a-bi)\\
=z^2+(-1-2i-a-bi)z+(1+2i)(a+bi)."


It follows that

"p+5i=-1-2i-a-bi\text{ and}\\
q(2-i)=(1+2i)(a+bi)=(a-2b)+(2a+b)i."

Therefore, we get the following system:


"\begin{cases}
p=-1-a\\
5=-2-b\\
2q=a-2b\\
-q=2a+b
\end{cases}"


which is equivalent to the following systems


"\begin{cases}
p=-1-a\\
b=-7\\
2q=a+14\\
-q=2a-7
\end{cases}"


| let us myltiply the last equation by 2 and add to the third equation |


"\begin{cases}
p=-1-a\\
b=-7\\
2q=a+14\\
0=5a
\end{cases}"


"\begin{cases}
p=-1\\
b=-7\\
q=7\\
a=0
\end{cases}"


We conclude that "p=-1,\ q=7" and the other root of the equation is "-7i."